OpenStudy (anonymous):

The wheel of a car is turning at rate 66 rev/min. Express the angular speed in rad/min. A little help please :)

OpenStudy (ash2326):

@kmeds16 Do you know how many radians the wheel turns in 1 revolution?

OpenStudy (anonymous):

2 pi? which is 360 degrees?

OpenStudy (ash2326):

yes, so we'll let it be 2 pi radians or 2*3.14= 6.28 rads per revolution

OpenStudy (anonymous):

uhmm aw okay.....

OpenStudy (ash2326):

How many revolutions the wheel is turning in a minute?

OpenStudy (anonymous):

66 rev/min

OpenStudy (ash2326):

So how many radian turn in a minute?

OpenStudy (anonymous):

do i have to multiply 66 and 2 pi? right?

OpenStudy (anonymous):

132?

OpenStudy (ash2326):

yes, but use 3.14 \[66\times 2\pi\] \[66\times 2 \times 3.14\] radians/min

OpenStudy (anonymous):

ahhhhh so it's 132pi rad/min :) got it. and i have to use 3.14... ok thanks!! :)

OpenStudy (ash2326):

welcome :)